OpenStudy (anonymous):

A series circuit has two 5.0-ohm resistors and a power source of 9.0 volts. How will the current be affected and what will it be if another resistor (with a resistance of 7.5 ohms) is added to the circuit? Select one of the options below as your answer: A. increase; 0.51 amps B. increase; 0.72 amps C. increase; 3.6 amps D. decrease; 0.51 amps E. decrease; 0.72 amps

OpenStudy (fifciol):

In situation before the another resistor is added: I=V/R=9/2*5=0,9 A In situation with added resistor: I=9/(2*5+7,5)=0,51, so it will decrease and get a finally value of 0.51 A Ans D

OpenStudy (anonymous):

wow your smart

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

DONT LAUGHT JPO FIFCIOL IS A GOD

OpenStudy (anonymous):

laugh* excuse my caps but this man fifciol is a legend pull game = 110%